# Profile-Aware Maneuvering: A Dynamic Multi-Agent System for Robust GAIA Problem Solving by AWorld

### v3

@@ 5 Analysis
To counteract this instability, we employ a dynamic maneuvering mechanism. Within this architecture, the Execution Agent is enabled to call upon a “Guard Agent” for review whenever it encounters a logical impasse. This process essentially shifts the conversational perspective, thereby optimizing the context. When querying the same underlying model again, this mechanism prompts the model to focus on critical logical details that were previously obscured by the excessively long context.
The Guard Agent generates more precise prompts as a refreshed context, helping to reorient the Execution Agent’s attention and facilitate convergence toward the correct answer. For instance, when tackling complex, grid-based constraint problems, the Guard Agent can identify and correct logical fallacies in the Execution Agent’s reasoning chain (a detailed case study is provided in a subsequent section). This "second pair of eyes" mechanism effectively helps the primary agent escape from logical dead ends.
@@
The limitations of reactive supervision highlight a central challenge in AI: how to effectively steer an agent’s behavior toward desired outcomes. The prevailing paradigm for this task is Reinforcement Learning (RL), which fine-tunes a model’s millions of opaque internal parameters based on scalar reward signals. This process implicitly shapes the agent’s decision-making policy, but the policy itself remains a "black box", a complex, distributed state within the model’s weights, making it difficult to interpret, audit, or directly control.
@@
The empirical evidence validates this philosophy. The dramatic results—a further 5.93% gain in pass@1 accuracy and a remarkable 57.41% reduction in standard deviation—demonstrate that this form of explicit, knowledge-based reinforcement is exceptionally effective. It proves that guiding an agent toward more optimal and stable behavior can be achieved more reliably by reinforcing the system with transparent, external knowledge, rather than solely attempting to tune the inscrutable depths of the model itself.
@@
Our framework, by contrast, constitutes a form of engineered introspection. The reinforcement mechanism—the performance fingerprint—is not a set of arcane weights but a human-readable text file. This transparency has profound implications for building dependable systems:
@@ 6 Future work
  • Towards Self-Aware Collectives: Online System Identification. A significant leap forward lies in enabling Online System Identification. This would allow the Guard Agent to dynamically update its behavioral fingerprint of the Execution Agent in real-time based on their ongoing interactions, learning from new successes and failures as they occur. This concept can be scaled to more complex, decentralized MAS, creating a "many-to-many" profiling architecture where each agent curates a dynamic portfolio of behavioral models for all its collaborators. Such a system would form a truly adaptive and self-aware collective, capable of optimizing its collaborative strategies on the fly.
  • The Symbiotic Assistant: Generalizing to Human-Agent Interaction. Perhaps the most ambitious extension is to generalize our paradigm beyond agent-agent interaction to the domain of human-agent collaboration. In this vision, an AI assistant would construct an empirical profile of its human user—their cognitive habits, common mistakes, or knowledge gaps. Armed with this deep understanding, the AI could offer truly personalized and proactive support, anticipating needs and preventing errors before they happen. Such research would elevate the "profile-aware" paradigm from a method for improving agent stability into a foundational engineering principle for designing the next generation of symbiotic intelligent systems.
@@ 7 Conclusion
In this work, we presented a novel framework for enhancing the robustness of Multi-Agent Systems by moving from generic supervision to profile-aware maneuvering. Our core contribution is the adaptation of System Identification, a principle from control theory, to create an automated methodology for modeling and understanding LLM agent behavior. We introduced a two-stage process within our AWorld platform: a preparatory offline stage to generate a "performance fingerprint" for our Execution Agent, and a crucial online stage where a Guard Agent leverages this fingerprint to provide targeted, preemptive guidance.
The practical efficacy of this approach was validated on the demanding GAIA benchmark. Our profile-aware MAS significantly outperformed both single-agent systems and its naive MAS counterpart, demonstrating marked improvements in effectiveness and stability. This superior performance culminated in our system achieving first place among open-source projects on the GAIA leaderboard. Ultimately, our findings suggest that the path toward truly resilient and trustworthy AI systems lies not just in promoting collaboration, but in equipping agents with a deep, data-driven understanding of their partners’ unique behavioral profiles.

### v4

@@ 5 Analysis
To counteract this instability, we employ a dynamic maneuvering mechanism. Within this architecture, the Execution Agent is enabled to call upon a Guard Agent for review whenever it encounters a logical impasse. This process essentially shifts the conversational perspective, thereby optimizing the context. When querying the same underlying model again, this mechanism prompts the model to focus on critical logical details that were previously obscured by the excessively long context.
The Guard Agent generates more precise prompts as a refreshed context, helping to reorient the Execution Agent’s attention and facilitate convergence toward the correct answer. For instance, when tackling complex, grid-based constraint problems, the Guard Agent can identify and correct logical fallacies in the Execution Agent’s reasoning chain (a detailed case study is provided in a subsequent section). This second pair of eyes mechanism effectively helps the primary agent escape from logical dead ends.
@@
The limitations of reactive supervision highlight a central challenge in AI: how to effectively steer an agent’s behavior toward desired outcomes. The prevailing paradigm for this task is Reinforcement Learning (RL), which fine-tunes a model’s millions of opaque internal parameters based on scalar reward signals. This process implicitly shapes the agent’s decision-making policy, but the policy itself remains a black box, a complex, distributed state within the model’s weights, making it difficult to interpret, audit, or directly control.
@@
The empirical evidence validates this philosophy. The dramatic results, a further 5.93% gain in Pass@1 accuracy and a remarkable 57.41% reduction in standard deviation, demonstrate that this form of explicit, knowledge-based reinforcement is exceptionally effective. It proves that guiding an agent toward more optimal and stable behavior can be achieved more reliably by reinforcing the system with transparent, external knowledge, rather than solely attempting to tune the inscrutable depths of the model itself.
@@
Our framework, by contrast, constitutes a form of engineered introspection. The reinforcement mechanism, the performance fingerprint, is not a set of arcane weights but a human-readable text file. This transparency has profound implications for building dependable systems:
@@ 6 Future work
  • Towards Self-Aware Collectives: Online System Identification. A significant leap forward lies in enabling Online System Identification. This would allow the Guard Agent to dynamically update its behavioral fingerprint of the Execution Agent in real-time based on their ongoing interactions, learning from new successes and failures as they occur. This concept can be scaled to more complex, decentralized MAS, creating a many-to-many profiling architecture where each agent curates a dynamic portfolio of behavioral models for all its collaborators. Such a system would form a truly adaptive and self-aware collective, capable of optimizing its collaborative strategies on the fly.
  • The Symbiotic Assistant: Generalizing to Human-Agent Interaction. Perhaps the most ambitious extension is to generalize our paradigm beyond agent-agent interaction to the domain of human-agent collaboration. In this vision, an AI assistant would construct an empirical profile of its human user, their cognitive habits, common mistakes, or knowledge gaps. Armed with this deep understanding, the AI could offer truly personalized and proactive support, anticipating needs and preventing errors before they happen. Such research would elevate the "profile-aware" paradigm from a method for improving agent stability into a foundational engineering principle for designing the next generation of symbiotic intelligent systems.
@@ 7 Conclusion
In this work, we introduce a new paradigm for engineering robust Multi-Agent Systems, moving beyond generic collaboration to the principled domain of predictive, model-based control of agent behavior. Our approach is grounded in Engineering Cybernetics, demonstrating that the principles used to maneuver complex physical systems like ships can be powerfully adapted to guide the reasoning trajectories of Large Language Model agents.
Our central contribution is the operationalization of this paradigm. We repurposed the classical technique of System Identification to create an automated, data-driven method for generating a performance fingerprint, an explicit model of an agent’s characteristic weaknesses. We then showed how this fingerprint transforms a standard Multi-Agent System, which acts as a simple reactive feedback loop, into a sophisticated Profile-Aware system that implements a composite feed-forward-feedback control strategy. This enables the supervisor agent to anticipate and preempt errors, rather than merely correcting them after the fact.
Validated on the demanding GAIA benchmark, our approach’s efficacy was demonstrated through a trifecta of empirical results. Our Profile-Aware MAS not only achieved superior overall accuracy, but also empirically validated the principles of robust control by: Taming Instability: Dramatically reducing the performance variance introduced by tool use by over 57%. Maximizing Reliability: Achieving the tightest gap between potential (Pass@3) and single-pass performance (Pass@1_avg), proving its ability to reliably convert its capability into correct actions.
This superior and more stable performance culminated in our system achieving first place among all open-source projects on the GAIA leaderboard at the time of submission. Ultimately, our findings suggest a pivotal shift in agent design methodology: from the empirical craft of prompt engineering towards the rigorous discipline of control systems engineering. The path toward truly resilient and trustworthy AI systems lies not just in promoting agent collaboration, but in building predictive models of their behavior and designing intelligent control architectures to maneuver them with precision and foresight.